If an airplane travel 600mph and a m-16 bullet travels at 2,300ft per sec, which one faster and by hom much?
Answer by mananth(16949)

600 mph
convert to feet / second
600 * 5280 /60*60 5280 feet = 1 mile : 60*60 seconds = 1 hour
=880 feet per second
2300-880=1420
The bullet is faster by 1420 feet/sec